# PIKA: Expert-Level Synthetic Datasets for Post-Training Alignment from Scratch

### v1

@@ 5 Related Work
Persona Roleplay. Persona-based roleplay offers a powerful mechanism for exploring the knowledge and perspectives embedded in LLMs. In this paradigm, the model’s world knowledge is effectively “compressed” into distributed representations, which can then be “decompressed” through the adoption of diverse personas to generate contextually grounded text (Delétang et al., 2024; Ge et al., 2024). Recent work by Ge et al. (2025) introduced the Persona Hub, an automatically constructed repository derived from large-scale web data. This resource allows researchers to tap into a wide variety of perspectives within LLMs, facilitating the large-scale creation of synthetic data without reliance on a small fixed seed corpus. Compared to traditional prompt-based synthesis, persona-driven approaches can be seamlessly integrated into arbitrary prompts, leveraging the strong role-playing capacity of LLMs to generate more diverse and versatile data. In our work, persona signals are randomly sampled from the Persona Hub to construct expert-level prompts.
Dataset Construction. Methods for building alignment datasets generally fall into two categories: human-in-the-loop collection and synthetic instruction generation. The first line of work gathers datasets through human interactions with LLMs (Databricks, 2023; Zhao et al., 2024; Zheng et al., 2024; 2023; Köpf et al., 2023). While such data is often high quality, it is expensive to scale, time-consuming to curate, and may contain toxic or biased content (Zhao et al., 2024). The second line of work leverages LLMs to generate synthetic instructions from a small set of human-annotated seeds, which are expanded using prompting strategies (Wang et al., 2023; Taori et al., 2023; Xu et al., 2023a; b; Wang et al., 2024; Sun et al., 2023). However, these methods often suffer from limited diversity, as generated instructions tend to resemble the original seeds (Li et al., 2024a). To address this, some approaches attempt to broaden coverage by leveraging summaries of world knowledge to drive generation (Ding et al., 2023; Li et al., 2024a). Our proposed PiKa dataset belongs to the synthetic category, but differs in that we employ persona-driven prompts sampled from expert-level profiles and selectively retain high-difficulty instances to better capture preference signals.
@@ 6 Conclusion
We present PiKa, a family of expert-level alignment datasets for LLMs. By leveraging persona-driven instruction generation and automatic filtering, we curated a compact yet diverse set of SFT and preference data that improves alignment quality. Fine-tuning experiments on Llama-3-8B show that models trained on PiKa-SFT outperform those trained on much larger open-source datasets and even surpass the official Llama-3-8B-Instruct model, trained on over 10M proprietary examples. We further validate PiKa’s effectiveness across different model families, such as Qwen2.5, consistently outperforming their official instruction-tuned counterparts. These results highlight PiKa’s potential to make alignment research more data-efficient and accessible to the open-source community, while suggesting future directions for studying data efficiency and mixture strategies in LLM alignment.
Appendix A Limitation
@@
Appendix B Ethics Statement
@@
Appendix C Evaluation Prompt
C.1 Difficulty evaluation prompt
@@
C.2 Realism evaluation prompt
@@
C.3 Quality evaluation prompt
@@
Appendix D Examples from PiKa Dataset
@@
D.1 Example 1: Complex Scientific Analysis
@@ Conclusion
D.2 Example 2: Medical Information Request
@@
Appendix E Use of Large Language Models in Paper Writing

### v2

@@ 5 Related Work
Persona Roleplay. Persona-based roleplay offers a powerful mechanism for exploring the knowledge and perspectives embedded in LLMs. In this paradigm, the model’s world knowledge is effectively “compressed” into distributed representations, which can then be “decompressed” through the adoption of diverse personas to generate contextually grounded text (delétang2024languagemodelingcompression; Ge et al., 2024). Recent work by Ge et al. (2025) introduced the Persona Hub, an automatically constructed repository derived from large-scale web data. This resource allows researchers to tap into a wide variety of perspectives within LLMs, facilitating the large-scale creation of synthetic data without reliance on a small fixed seed corpus. Compared to traditional prompt-based synthesis, persona-driven approaches can be seamlessly integrated into arbitrary prompts, leveraging the strong role-playing capacity of LLMs to generate more diverse and versatile data. In our work, persona signals are randomly sampled from the Persona Hub to construct expert-level prompts.
Dataset Construction. Methods for building alignment datasets generally fall into two categories: human-in-the-loop collection and synthetic instruction generation. The first line of work gathers datasets through human interactions with LLMs (Databricks, 2023; Zhao et al., 2024; Zheng et al., 2024, 2023; Köpf et al., 2023). While such data is often high quality, it is expensive to scale, time-consuming to curate, and may contain toxic or biased content (Zhao et al., 2024). The second line of work leverages LLMs to generate synthetic instructions from a small set of human-annotated seeds, which are expanded using prompting strategies (Wang et al., 2023; Taori et al., 2023; Xu et al., 2023a, b; Wang et al., 2024; Sun et al., 2023). However, these methods often suffer from limited diversity, as generated instructions tend to resemble the original seeds (Li et al., 2024a). To address this, some approaches attempt to broaden coverage by leveraging summaries of world knowledge to drive generation (Ding et al., 2023; Li et al., 2024a). Our proposed PiKa dataset belongs to the synthetic category, but differs in that we employ persona-driven prompts sampled from expert-level profiles and selectively retain high-difficulty instances to better capture preference signals.
@@ 6 Conclusion
We present PiKa, a family of expert-level alignment datasets for LLMs. By leveraging persona-driven instruction generation and automatic filtering, we curated a compact yet diverse set of SFT and preference data that improves alignment quality. Fine-tuning experiments on Llama-3-8B show that models trained on PiKa-SFT outperform those trained on much larger open-source datasets and even surpass the official Llama-3-8B-Instruct model, trained on over 10M proprietary examples. We further validate PiKa’s effectiveness across different model families, consistently outperforming their official instruction-tuned counterparts. These results highlight PiKa’s potential to make alignment research more data-efficient for the open-source community.
7 Limitation
@@
Appendix A Ethics Statement
@@
Appendix B Evaluation Prompt
B.1 Difficulty evaluation prompt
@@
B.2 Realism evaluation prompt
@@
B.3 Quality evaluation prompt
@@
Appendix C Examples from PiKa Dataset
@@
C.1 Example 1: Complex Scientific Analysis
@@ Conclusion
C.2 Example 2: Medical Information Request
@@
Appendix D Use of Large Language Models in Paper Writing
@@
Appendix E Additional Analyses and Clarifications
E.1 Effect of Prompt Difficulty
PiKa adopts a difficulty-driven construction strategy. To examine whether prompt difficulty causally improves alignment performance, we apply a down-quality reconstruction procedure to PiKa prompts and regenerate lower-difficulty variants using the guided template shown below.
System Prompt: You are an expert evaluator for assessing the difficulty level of instructions/questions. Please evaluate the given instruction on a scale of 1–10 based on: Complexity of the task or question Level of domain knowledge required Cognitive load needed to understand and process Technical depth and sophistication Scoring guidelines: 1–2: Very basic, elementary level knowledge 3–4: Intermediate, requires some specialized knowledge 5–6: Advanced, requires solid domain expertise 7–8: Expert level, requires deep specialized knowledge 9–10: Cutting-edge, requires highly specialized or research-level expertise Now, instead of giving a score, please rewrite the given instruction into a new version that would have a difficulty score of 1–2 according to the above scale. Guidelines: Keep the topic and domain consistent Simplify the reasoning and remove complexity Avoid multi-step or abstract reasoning Output strictly in JSON: { ”instruction”: ”¡rewritten easy instruction¿” }
As shown in Table 3, PiKa exhibits substantially higher prompt difficulty than Magpie (average 7.39 vs. 2.65). To more directly test the causal effect of difficulty, we reconstruct lower-quality prompts and regenerate responses using the above template. Each level reduces prompt difficulty while holding other factors fixed. The “mid” variant is derived from the vanilla prompt, and the “low” variant is derived from the “mid” prompt. Difficulty is re-evaluated with GPT-4o.
$\begin{split}P^{\prime}_{\text{new}}&=\mathrm{LLM}(P_{\text{vanilla}},\text{score\_guide}),\\ R_{\text{new}}&=\mathrm{LLM}(P^{\prime}_{\text{new}}).\end{split}$ (7)
The results in Table 3 confirm that higher prompt difficulty leads to stronger alignment gradients.
Figure 6: Example instructions from Magpie-Pro and PiKa datasets. PiKa instructions are generally more complex and domain-specific, while Magpie-Pro tends to focus on simpler, open-ended prompts. We also include detailed instruction-response pairs in Appendix C.
E.2 Effect of Persona Conditioning
Persona conditioning encourages generation of more specialized, higher-difficulty, and less redundant prompts. Removing persona information substantially reduces both difficulty and diversity. Results are shown in Table 6.
Table 6: Persona ablation (10k samples). Mean MND (Minimum Neighbor Distance) measures instruction redundancy (higher is better).
Model (10k) Persona Difficulty Mean MND LC (%) WR (%)
LLaMA-3-8B-PiKa none 3.11 0.0234 13.84 15.53
LLaMA-3-8B-PiKa enabled 7.39 0.497 31.01 30.32
E.3 Persona Domain Distribution
We analyze persona-domain coverage across the PiKa-30k dataset. Table 7 shows that personas largely fall within broad, widely applicable domains rather than narrow or niche specialties.
Table 7: Top 10 persona-domain distribution in PiKa (percentage).
Domain Percentage
General 36.3%
Medicine 16.0%
Biology 13.3%
History 12.5%
Engineering 6.8%
Environmental Science 6.5%
Education 5.8%
Psychology 4.3%
Computer Science 4.0%
To assess robustness to domain composition, we train three PiKa-10k subsets sampled with different seeds. Variance across runs is minimal.
Table 8: Robustness across different random samples of PiKa (10k).
Experiment LC (%) WR (%)
1 31.01 30.32
2 30.09 30.17
3 30.45 30.06
E.4 Evaluation Bias Across Different Judges
To examine whether using GPT-4o as an evaluator introduces systematic bias, we further evaluate PiKa, LLaMA-3-Instruct, and Magpie-Pro using two independent judges: GPT-5 and GPT-4.1-mini. Results in Table 9 show that PiKa consistently outperforms baselines across all evaluators, indicating robustness to choice of judge.
Table 9: Evaluation across different judges.
Judge Model LC (%) WR (%) SD N Avg Len
GPT-5 LLaMA-3-8B-PiKa 19.28 19.00 1.35 805 1927
LLaMA-3-8B-Instruct 15.95 16.56 1.28 805 1949
LLaMA-3-8B-Magpie-Pro 9.51 14.55 1.19 805 2461
GPT-4.1-mini LLaMA-3-8B-PiKa 29.02 27.55 1.52 805 1927
LLaMA-3-8B-Instruct 25.44 25.53 1.45 805 1949
LLaMA-3-8B-Magpie-Pro 20.17 28.54 1.53 805 2461
E.5 Length Analysis Across 5,000 Prompts
We analyze 5,000 prompts, each paired with five candidate responses, to examine whether verbosity bias affects reward-model preferences. Reported lengths correspond to character lengths (approximately 1,300 tokens). Across all prompts, the chosen responses are on average only 7.5% longer than the rejected ones (5,343 vs. 4,969 characters). Moreover, only 30.8% of prompts select the longest candidate as the preferred response, and the global correlation between response length and reward score remains mild ($r=0.31$). These observations collectively indicate that length is not the dominant factor driving reward-model decisions.
Table 10: Length analysis across 5,000 prompts.
Metric Chosen Response Rejected Response $\Delta\%$ Longest Selected (%) Corr (r)
Length (chars) 5,343 4,969 +7.5% 30.8% 0.31
E.6 Robustness Across Different Reward Models
To examine whether PiKa’s evaluation results depend on a particular reward model, we additionally compare two open-source reward models in the 10K setting: Skywork-Reward-V2-LLaMA-3.1-8B and ArmoRM-Llama3-8B-v0.1. As shown in Table 11, the resulting AlpacaEval 2 scores are highly consistent across the two choices. Skywork-Reward-V2-LLaMA-3.1-8B achieves 31.01 LC and 30.32 WR, while ArmoRM-Llama3-8B-v0.1 obtains 29.85 LC and 30.05 WR. The small performance gap suggests that PiKa’s gains are not tied to a single evaluator and remain stable across different open-source reward models.
Reward Model LC WR
Skywork-Reward-V2-LLaMA-3.1-8B 31.01 30.32
RLHFlow/ArmoRM-Llama3-8B-v0.1 29.85 30.05
Table 11: AlpacaEval 2 results under different open-source reward models in the 10K setting.
E.7 Robustness Across Multiple LLM Judges
To assess whether the data-quality conclusions are sensitive to the choice of judge, we further evaluate PiKa and Magpie-Pro using multiple LLM judges, including GPT-4o, Gemini-2.5-Flash, and Gemini-3-Flash. As shown in Table 12, the same overall trend holds across all three judges: PiKa is consistently rated as substantially more difficult than Magpie-Pro, while also achieving slightly higher quality scores on average. These results suggest that the data-quality comparison remains robust across different judges.
Dataset GPT-4o Gemini-2.5-Flash Gemini-3-Flash Avg.
Difficulty score
PiKa 7.39 7.98 6.45 7.27
Magpie-Pro 2.65 4.51 3.17 3.44
Quality score
PiKa 9.57 9.53 9.19 9.43
Magpie-Pro 9.01 9.37 9.03 9.14
Table 12: Data-quality evaluation results across multiple LLM judges. PiKa is consistently rated as more difficult than Magpie-Pro, while also achieving slightly higher quality scores on average.
E.8 Comparison with Difficulty-Matched Filtering
A natural question is whether PiKa’s gains can be recovered simply by selecting the highest-scoring subset from an existing synthetic dataset. To examine this, we construct a control set by selecting the top 30K examples from Magpie-Pro using the same Skywork-Reward-V2 reward model. As shown in Table 13, this filtered subset still has a much lower average difficulty than PiKa-30K (3.01 vs. 7.39), indicating that it does not match the higher-difficulty prompt distribution targeted by PiKa. Correspondingly, PiKa-30K still achieves substantially better AlpacaEval 2 LC performance than the filtered Magpie-Pro subset (32.82 vs. 20.86). These results suggest that PiKa’s advantage cannot be explained solely by reward-model filtering over an existing dataset, but instead comes from producing a substantially harder and more effective training distribution.
Dataset Avg. Difficulty LC
Magpie-Pro-Top-30K 3.01 20.86
PiKa-30K 7.39 32.82
Table 13: Comparison between PiKa-30K and a reward-filtered top-30K subset from Magpie-Pro selected by the same Skywork-Reward-V2 model.
E.9 Controlled Comparison with Different Generators
To further isolate the role of the generator, we conduct a controlled comparison under the same expert-level prompt setting by replacing the GPT-4o generator in PiKa with Llama-3-70B-Instruct, which is also used in Magpie. As shown in Table 14, PiKa with Llama-3-70B-Instruct preserves the high prompt difficulty of PiKa, matching PiKa with GPT-4o on difficulty (7.39 vs. 7.39). However, its response quality is substantially lower than PiKa with GPT-4o (7.01 vs. 9.57), and its downstream AlpacaEval 2 performance also drops markedly. In particular, PiKa with Llama-3-70B-Instruct achieves only 14.59 LC and 16.27 WR, compared with 31.01 LC and 30.32 WR for PiKa with GPT-4o. These results suggest that under expert-level, high-difficulty prompts, a weaker generator cannot produce responses of comparable quality, which further supports the need for a stronger GPT-based generator in our pipeline.
Dataset Difficulty Response Quality LC (%) WR (%)
Magpie-Pro 2.65 9.01 15.42 16.89
PiKa w/ Llama-3-70B-Instruct 7.39 7.01 14.59 16.27
PiKa w/ GPT-4o 7.39 9.57 31.01 30.32
Table 14: Controlled comparison under the same expert-level prompt setting with different generators. Replacing GPT-4o with Llama-3-70B-Instruct preserves prompt difficulty but substantially reduces response quality and downstream chat performance.